OpenStudy (anonymous):

Factor completely 5x^2 + 30x + 50 5(x + 2)(x + 5) 5(x^2 + 6x + 10) 5x(x + 6 + 10) 5x(x + 2)(x + 5)

OpenStudy (anonymous):

\[5x^2 + 30x + 50=5(x^2+6x+10)\]

OpenStudy (anonymous):

D=6*6-4*10=36-40=-6 it means you can't factor it more

OpenStudy (anonymous):

5(x^2 + 6x + 10)

OpenStudy (anonymous):

5(x^2+6x+10)